An 11.25 g ethanol is heated from 25 degree C to 52.1 degree C. Calculate the amount of heat absorbed by the ethanol.


1
Expert's answer
2021-04-27T07:48:42-0400

According to the heat equation:

Q = cm(T2-T1)

where Q - absorbed heat, c - specific heat (2.44 J/gºC), m - mass, T2 - final temperature, T1 - initial temperature.

From here:

Q(ethanol) = cm(T2-T1) = 2.44 J/gºC × 11.25 g × (52.1ºC - 25ºC) = 744 J


Answer: 744 J